OpenStudy (konradzuse):

solve for x ...... -5 + 7e^3x = 10

OpenStudy (konradzuse):

+5 to each side yeilds 7e^(3x) = 15 /7 on each side e^(3x) = 15/7 lne^(3x) = ln(15/7) 3x = ln(15/7) x = ln(15/7) / 3?

OpenStudy (anonymous):

Looks good!